OpenStudy (anonymous):

Help with math problem? The directions say to solve the equation and then check your solution. 5/9d=9/10

OpenStudy (darkbluechocobo):

I believe you would Cross multiply to find the answer: 81d=50

OpenStudy (darkbluechocobo):

Mehh I am not sure /.\

OpenStudy (darkbluechocobo):

because if you did that it would come out 0.6172839506172

OpenStudy (darkbluechocobo):

I found the answer :p 5/9 d = 9/10. d = (9/10) / (5/9) = 1.62. d = 1.62.